OpenStudy (anonymous):

Let f(x) = x + 9 and g(x) = x – 3. What is f(x) • g(x)

OpenStudy (anonymous):

You use FOIL: \[ \Large{(x+9) \cdot (x-3) = x^2 + 9x - 3x - 27 = \boxed {x^2 + 6x - 27}.} \]

OpenStudy (anonymous):

oh thank you that is the answer

OpenStudy (anonymous):

No problem! Want to give me a medal? =P Just saying ...

OpenStudy (anonymous):

yes how to do tht

OpenStudy (anonymous):

Click "Best Response" on one of my posts. Oh and thanks for becoming my fan! :)